OpenStudy (hanner_b_nanner):

Fan&Medal...Can someone check my answer please? An observer (O) is located 500 feet from a school (S). The observer notices a bird (B) flying at a 39° angle of elevation from his line of sight. What equation and trigonometric function can be used to solve for the height (h) of the bird? What is the height of the bird? Pick another trigonometric function and describe why that function is not appropriate when trying to solve for (h). You must show all work and calculations to receive full credit.

OpenStudy (hanner_b_nanner):

I was really confused...but I got this. We need to find the sine. sin=opposite/hypotenuse sin 39=x/500 sin 39(10)=x/10(10) sin 39(10)=x sin 39(10)= 6.29 If you were to use the tangent, you wouldn't be able to come to this exact number. You would only have a simplified answer, and not the exact.

OpenStudy (cj49):

We know the side adjacent to the angle and we want to know the opposite side.

OpenStudy (cj49):

tan=opp/hypo